# Security of Language Models for Code: A Systematic Literature Review

### v1

@@ 3. Survey Methodology
In this paper, we follow the systematic literature review (SLR) methodology proposed by Kitchenham et al. (Staffs et al., 2007; Kitchenham et al., 2023) to conduct this survey. This method is widely utilized in the majority of other SLRs related to SE (Ramírez et al., 2019; Liu et al., 2023; Wang et al., 2023b; Zhang et al., 2024a; Hou et al., 2023). Our methodology consists of three main steps: planning the review, conducting the review, and analyzing the review results.
@@ 3.3. Search Strategy
Our search strategy integrates manual and automated methods, employing high-quality search strings and comprehensive databases to avoid omissions. Initially, we conduct manual searches to identify a set of relevant studies. From these studies, we derive a search string. Next, we perform automated searches across six commonly used databases to maximize coverage and minimize the risk of overlooking relevant research. Finally, we implement a series of relatively stringent filtering steps to obtain relevant studies.
@@ 3.3.1. Search Strings
Firstly, to obtain suitable search strings, we follow (Zhang et al., 2011) and conduct a manual search to identify relevant conferences. Our research topic, “security of language models for code” involves both SE and model security. Therefore, we select six top conferences and journals in SE and six in network and information security, as shown in Table 1, for manual searches of papers related to CodeLM security. Initially, we systematically crawl a list of top academic publications containing 10,267 papers. Subsequently, we use scripts for automated scanning and manual verification to confirm 14 papers relevant to our research objectives. Next, we extract the search string from those papers. The search string should combine two sets of keywords: one related to CodeLMs and the other related to model security. To avoid omitting papers related to our research as much as possible, we follow (Hou et al., 2023) and expand the keywords related to CodeLMs by incorporating terms relevant to LMs and additional tasks related to code. Only when a paper includes both sets of keywords is it more likely to be the paper we need. The complete set of search keywords is as follows:
@@ 3.3.2. Search Sources
After determining the search string through the manual search, we conduct automated searches across six widely used databases (namely, Google Scholar, ACM Digital Library, IEEE Xplore, Springer, DBLP, and ArXiv), which are standard and common online engines or databases (of the Planet of Serverless Computing: A Systematic Review, 2023; Sun et al., 2024d; Zhang et al., 2024a). It is worth noting that, considering that some researchers are inclined to disclose the latest technique on ArXiv in advance, we included ArXiv in our search databases. Considering that in 2018, Meng et al. (Meng et al., 2018) was the first to focus on the security of CodeLMs, our search spanned from 2018 to August 2024. Firstly, we used combinations of two sets of search keywords (a total of 407 combinations) for precise searches in each database, and the results were merged and deduplicated. Specifically, we obtained 7,484 papers from Google Scholar, 1,247 papers from ACM Digital Library, 257 papers from IEEE Xplore, 4,348 papers from Springer, 302 papers from DBLP, and 914 papers from ArXiv.
@@ 3.3.3. Study Selection
Figure 2. Study identification and selection process.
[…]
Study identification and selection process.
@@
In the first phase, we conduct automatic filtering based on C1 to exclude papers not written in English, reducing the number of papers to 12,071. Next, we retain only journal and conference papers, filtering out books and theses (meeting C2). We use scripts to automatically extract the BibTeX of the candidate papers. The first part of the BibTeX entry indicates the type of publication. Books are labeled as @book, @booklet, or @inbook, while Master’s theses are denoted as @masterthesis, and Ph.D. theses as @phdthesis. Journal papers are classified as @article, and conference papers as @inproceedings or @conference. Therefore, based on the information of the first part of the BibTeX entry, we use automated scripts to only include entries categorized as @article for journal papers and @inproceedings or @conference for conference papers. After filtering, we obtain 8,851 papers. Furthermore, we filter the papers manually based on their titles, abstracts, keywords, or full texts. In the third phase, we evaluate the remaining papers based on their titles and abstracts to determine if they meet the requirements of C3. We filter out technical reports, literature reviews and surveys, leaving only 7,218 technical papers. In the subsequent phases, we assess whether the papers contained at least one CodeLM by examining their full texts, i.e., whether they meet the requirements of C4. We identify 461 papers that meet the criteria. This phase aims to narrow down the scope and directly identify papers relevant to CodeLMs. Finally, we review the full texts of the remaining papers once more and selected only those that explicitly mention at least one approach aimed at CodeLM security (meeting C5). This phase primarily excluded papers solely focused on CodeLM performance validation without addressing CodeLM security. We ultimately select 67 papers that meet the criteria.
@@ 3.4. Snowballing Search
We consider the 60 papers after the study selection process as the initial paper list. Subsequently, we iteratively repeat the snowballing process until no new relevant papers are discovered. We perform backward snowballing and forward snowballing to collect 476 and 1,242 papers respectively. After deduplication, including deduplication of the initial paper list, 1,645 papers still remain. Then, we conduct a comprehensive study selection process on these 1,645 papers, including filtering criteria C1 to C5. Finally, we obtain an additional 7 papers.
@@ 3.5. Collection Results
ISSTA International Symposium on Software Testing and Analysis 2024 (Sun et al., 2024b)
@@
TOSEM ACM Transactions on Software Engineering and Methodology 2022 2024 (Zhang et al., 2022a; Zhou et al., 2022) (Li et al., 2024b; Yao et al., 2024)
@@
ACL Annual Meeting of the Association for Computational Linguistics 2023 (Sun et al., 2023a; Li et al., 2023b; Na et al., 2023)
@@
EMNLP Conference on Empirical Methods in Natural Language Processing 2022 (Choi et al., 2022)
@@
USENIX Security USENIX Security Symposium 2019 2021 2023 (Quiring et al., 2019) (Schuster et al., 2021; Severi et al., 2021) (Zhang et al., 2023d)
@@
FGCS Future Generation Computer Systems 2023 (Li et al., 2023a)
@@
CollaborateCom International Conference on Collaborative Computing: Networking, Applications and Worksharing 2024 (Zhang et al., 2024d)
@@
ArXiv The arXiv. org e-Print archive 2018 2022 2023 2024 (Meng et al., 2018) (Jia et al., 2022) (Wu et al., 2023; Hussain et al., 2023a; Zhang et al., 2023c; Qi et al., 2023) (Zhang et al., 2023e; Abuhamad et al., 2023; Improta et al., 2023; Nguyen et al., 2023; Capozzi et al., 2023) (Hussain et al., 2024a, b; Yang et al., 2024d; Awal et al., 2024)
Figure 3. Accumulation and venue distribution of publications related to security of CodeLMs from 2018 to August 2024.
@@
Table 2 lists the detailed publication venues and years of the 67 reviewed papers. An overview of the distribution of included papers is shown in Figure 3. Figure 3 (a) shows the cumulative distribution of publications included from 2018 to August 2024. In 2018, there are only 1 relevant paper. Since 2021, the number of relevant papers published each year has remained consistently at or above 10. By 2023, the total number of publications in this field have reached 51. As of August 2024, the number of publications in this field has reached 67. This rapid growth trend demonstrates a growing research interest in the security domain of CodeLMs, and the importance of security attributes for CodeLMs is increasing.
Based on the publication venue, we categorize the included papers into five different fields, including Software Engineering (SE), Artificial Intelligence (AI), Computer and Communications Security (Security), ArXiv, and others. Figure 3 (b) shows the distribution of publications published in different research fields. Among all the papers, 77.6% are published in peer-reviewed conferences or journals. 35.8% of the papers are published in software engineering venues, such as ICSE, ESEC/FSE, and so on. 14.9% are published in security venues, and another 14.9% in AI venues. Additionally, 11.9% of the papers are published in other fields. This indicates that the importance of CodeLM security has gained attention across multiple fields. Additionally, the remaining 22.4% of the papers have not yet been published in peer-reviewed venues and are instead available on ArXiv, an open-access academic platform.
@@ 4. Answer to RQ1: Attacks on CodeLMs
As shown in the left part of Figure 4, we construct a taxonomy based on the 53 research papers collected on attacks against CodeLMs. Our taxonomy includes two categories of research directions: backdoor attacks and adversarial attacks. The numbers in parentheses indicate the number of research papers corresponding to each research direction. For example, there are 13 papers that focus on studying backdoor attacks in CodeLMs and 40 papers that are dedicated to researching adversarial attacks in CodeLMs. Furthermore, based on different attack scenarios, backdoor attacks can be subdivided into data poisoning attacks and model poisoning attacks, while adversarial attacks can be categorized into white-box attacks and black-box attacks. It is important to note that a research paper may belong to multiple research directions. For example, Schuster et al. (Schuster et al., 2021) proposes both data poisoning and model poisoning techniques. In backdoor attacks, research on data poisoning attacks and model poisoning attacks account for 76.92% (10/13) and 30.77% (4/13), respectively. It can be observed that existing research focuses more on data poisoning attacks. In adversarial attacks, research on white-box attacks and black-box attacks account for 20% (8/40) and 80% (32/40), respectively, with black-box attacks receiving greater attention.
@@ 4.2.1. Overview of Backdoor Attacks against CodeLMs
Figure 5 shows the general workflow of backdoor attacks against CodeLMs. Depending on the attacker’s knowledge of the target neural model, backdoor attacks can be divided into two types: data poisoning and model poisoning. The workflow of the data poisoning is shown in Figure 5. Data poisoning aims to inject a specific trigger into input samples, causing models trained on these data to misclassify any input containing the trigger as the target label (Gu et al., 2017). Given clean samples $D=\{\mathcal{X},\mathcal{Y}\}$, where $x=\{x_{i}\}^{n}_{i=1}\in\mathcal{X}$ is a sequence of $n$ tokens. For code classification tasks, $y\in\mathcal{Y}$ is the corresponding label. The attacker aims to generate a secret trigger with $m$ tokens $t^{*}=\{t^{*}_{i=1}\}^{m}_{i=1}$ and construct a set of poisoned samples $\mathcal{D}_{p}=\{\mathcal{D}\cup\mathcal{D}^{*}\}$, where $\mathcal{D}^{*}=\{\mathcal{X}^{*},y^{*}\}$, $x^{*}=\{x_{i}\}^{n}_{i=1}\oplus\{t^{*}_{i}\}^{m}_{i=1}\in\mathcal{X}^{*}$. $\oplus$ denotes the trigger injection operation. For code generation tasks, the corresponding ground truth can be represented as $\{y_{1},y_{2},\ldots,y_{n}\}\in\mathcal{Y}$, and the target label can be represented as $y^{*}=\{y_{1},y_{2},\ldots,y_{t},\ldots,y_{n}\}$, where the target label $y^{*}$ is generated by inserting the target token $y_{t}$ into the corresponding label. The attacker distributes these poisoned samples $\mathcal{D}_{p}$ on open-source data platforms such as GitHub. Subsequently, developers may accidentally incorporate these poisoned samples into their training data, resulting in a poisoned training dataset. This poisoned training dataset is then used to train the CodeLMs. During this process, a backdoor is stealthily implanted into the CodeLMs. The attacker can launch a backdoor attack using the pre-defined trigger $t^{*}$, causing the CodeLMs to output the target prediction.
The workflow of the model poisoning is also shown in Figure 5. Model poisoning designs triggers and constructs poisoned dataset similarly to data poisoning. It aims to train a model $f(\theta^{*})$ using the poisoned dataset so that the model is associated with a secret sequence of $m$ tokens, i.e., the trigger $t^{*}$ and a target label $y^{*}$. Model poisoning minimizes the following training loss.
@@
4.2.2. Data Poisoning Attacks.
@@
Ramakrishnan et al. (Ramakrishnan and Albarghouthi, 2022) propose a simple yet effective poisoning method targeting CodeLMs, which utilizes fixed or grammar dead code snippets as triggers and static or dynamic labels as attack targets. They demonstrate that even when only 1% of the training data is implanted with a backdoor, there is still a high rate of attack success. They conduct experiments using Code2Seq (Alon et al., 2019a) and Seq2Seq (Bahdanau et al., 2015) models, focusing on code summarization tasks and method name prediction tasks in Java and Python programming languages.
Schuster et al. (Schuster et al., 2021) explore data poisoning attacks on two advanced CodeLMs for code completion, Pythia (Svyatkovskiy et al., 2019) and GPT-2 (Radford et al., 2019). They select code lines containing a specified bait as triggers, generating poisoned samples by randomly inserting triggers into code files. They introduce a novel targeted attack that can impact only specific users. Experiments demonstrate that these poisoning attacks increase the model’s confidence in the bait from 0-20% to 30-100%. Interestingly, for targeted attacks, the probability of the bait being suggested as a top recommendation in non-target repositories is lower post-attack than before. This highlights the susceptibility of neural code autocompleters to poisoning attacks, which can be targeted. They also show that two defense methods, spectrum-based methods (Tran et al., 2018) and clustering methods (Chen et al., 2019b), incorrectly filter out most of the legitimate corpus content, retaining many poisoned files from attackers, resulting in a high false positive rate and ineffective defense.
Severi et al. (Severi et al., 2021) propose a universal, model-agnostic approach for backdoor attacks, aiming to investigate the sensitivity of ML-based malware classifiers to data poisoning attacks. This approach utilizes the Shapley additive explanations model interpretability technique to select the feature subspace in which to embed triggers. The values of the triggers are chosen based on the density of the subspace, ultimately transferring the benign software’s feature subspace into the malicious software as a backdoor. Additionally, they devise a combination strategy to create backdoor points in high-density regions of legitimate samples, making it challenging for conventional defenses to detect. They conduct attacks on Windows PE files, Android, and PDF files using the EMBER (Anderson and Roth, 2018), Contagio (Smutz and Stavrou, 2012), and Drebin (Arp et al., 2014) datasets. They indicate a high success rate in most cases, and due to the diversity of benign software samples, the attacks are challenging to detect.
Wan et al. (Wan et al., 2022) also utilize fixed or grammar dead code snippets as triggers, injecting them into natural language queries-code data containing the target word to create poisoned samples. They
@@
introduce data poisoning into CodeLMs for code search tasks, emphasizing the ongoing need for attention to the robustness and security of such models. Their experimental results demonstrate the effectiveness of this attack on CodeLMs based on BiRNN (Cho et al., 2014), Transformer, and CodeBERT, with simpler models exhibiting more successful attacks.
@@
Sun et al. (Sun et al., 2023a) indicate that the dead code-based triggers are low stealthiness and are easily detected. They propose a backdoor attack method named BADCODE targeting CodeLMs for code search tasks, to address the problem of low stealthiness of the trigger. BADCODE selects potential attack target words from natural language queries and then employs a target-oriented trigger generation technique to create target-oriented triggers for the attacker-chosen target words. Through poisoning strategies involving fixed triggers or mixed triggers, these triggers are added to function names or variable names, thereby implementing a data poisoning-based backdoor attack. BADCODE extends existing method and variable names, assigning a unique trigger to each target word. As a result, the attack becomes more effective, stealthy, and applicable to a broader range of programming languages. Experimental results indicate that BADCODE outperforms fixed trigger attacks and grammar trigger attacks (Wan et al., 2022)by 60% in backdoor attacks on code search LMs based on CodeBERT and CodeT5 (Wang et al., 2021a). Additionally, its stealthiness is twice as effective as the fixed trigger attacks and grammar trigger attacks (Wan et al., 2022).

### v2

@@ 3. Survey Methodology
In this paper, we follow the systematic literature review (SLR) methodology proposed by Kitchenham et al. (Staffs et al., 2007; Kitchenham et al., 2023) to conduct this survey. This method is widely utilized in the majority of other SLRs related to SE (Ramírez et al., 2019; Liu et al., 2023; Wang et al., 2023b; Zhang et al., 2024a; Hou et al., 2024). Our methodology consists of three main steps: planning the review, conducting the review, and analyzing the review results.
@@ 3.3. Search Strategy
Figure 2. Study identification and selection process.
[…]
Study identification and selection process.
Our search strategy integrates manual and automated methods, employing high-quality search strings and comprehensive databases to avoid omissions. Initially, we conduct manual searches to identify a set of relevant studies. From these studies, we derive a search string. Next, we perform automated searches across six commonly used databases to maximize coverage and minimize the risk of overlooking relevant research. Finally, we implement a series of relatively stringent filtering steps to obtain relevant studies. Figure 2 illustrates the process of study identification and selection.
@@ 3.3.1. Search Strings
Firstly, to obtain suitable search strings, we follow (Zhang et al., 2011) and conduct a manual search to identify relevant conferences. Our research topic, “security of language models for code” involves both SE and model security. Therefore, we select six top conferences and journals in SE and six in network and information security, as shown in Table 2, for manual searches of papers related to CodeLM security. Initially, we systematically crawl a list of top academic publications containing 10,267 papers. Subsequently, we use scripts for automated scanning and manual verification to confirm 14 papers relevant to our research objectives. Next, we extract the search string from those papers. The search string should combine two sets of keywords: one related to CodeLMs and the other related to model security. To avoid omitting papers related to our research as much as possible, we follow (Hou et al., 2024) and expand the keywords related to CodeLMs by incorporating terms relevant to LMs and additional tasks related to code. Only when a paper includes both sets of keywords is it more likely to be the paper we need. The complete set of search keywords is as follows:
@@ 3.3.2. Search Sources
After determining the search string through the manual search, we conduct automated searches across six widely used databases (namely, Google Scholar, ACM Digital Library, IEEE Xplore, Springer, DBLP, and ArXiv), which are standard and common online engines or databases (of the Planet of Serverless Computing: A Systematic Review, 2023; Sun et al., 2024b; Zhang et al., 2024a). It is worth noting that, considering that some researchers are inclined to disclose the latest technique on ArXiv in advance, we included ArXiv in our search databases. Considering that in 2018, Meng et al. (Meng et al., 2018) was the first to focus on the security of CodeLMs, our search spanned from 2018 to August 2024. Firstly, we used combinations of two sets of search keywords (a total of 407 combinations) for precise searches in each database, and the results were merged and deduplicated. Specifically, we obtained 7,484 papers from Google Scholar, 1,247 papers from ACM Digital Library, 257 papers from IEEE Xplore, 4,348 papers from Springer, 302 papers from DBLP, and 914 papers from ArXiv.
@@ 3.3.3. Study Selection
In the first phase, we conduct automatic filtering based on C1 to exclude papers not written in English, reducing the number of papers to 12,071. Next, we retain only journal and conference papers, filtering out books and theses (meeting C2). We use scripts to automatically extract the BibTeX of the candidate papers. The first part of the BibTeX entry indicates the type of publication. Books are labeled as @book, @booklet, or @inbook, while Master’s theses are denoted as @masterthesis, and Ph.D. theses as @phdthesis. Journal papers are classified as @article, and conference papers as @inproceedings or @conference. Therefore, based on the information of the first part of the BibTeX entry, we use automated scripts to only include entries categorized as @article for journal papers and @inproceedings or @conference for conference papers. After filtering, we obtain 8,851 papers. Furthermore, we filter the papers manually based on their titles, abstracts, keywords, or full texts. In the third phase, we evaluate the remaining papers based on their titles and abstracts to determine if they meet the requirements of C3. We filter out technical reports, literature reviews and surveys, leaving only 7,218 technical papers. In the subsequent phases, we assess whether the papers contained at least one CodeLM by examining their full texts, i.e., whether they meet the requirements of C4. We identify 461 papers that meet the criteria. This phase aims to narrow down the scope and directly identify papers relevant to CodeLMs. Finally, we review the full texts of the remaining papers once more and selected only those that explicitly mention at least one approach aimed at CodeLM security (meeting C5). This phase primarily excluded papers solely focused on CodeLM performance validation without addressing CodeLM security. We ultimately select 60 papers that meet the criteria.
@@ 3.4. Snowballing Search
We consider the 60 papers after the study selection process as the initial paper list. Subsequently, we iteratively repeat the snowballing process until no new relevant papers are discovered. We perform backward snowballing and forward snowballing to collect 476 and 1,242 papers respectively. After deduplication, including deduplication of the initial paper list, 1,645 papers still remain. Then, we conduct a comprehensive study selection process on these 1,645 papers, including filtering criteria C1 to C5. Finally, we obtain an additional 8 papers. For the final set of 68 papers, the first two authors review them thoroughly and multiple times to ensure that each paper has a clear research motivation, provides detailed descriptions of the techniques, presents comprehensive experimental setups, and clearly confirms the experimental findings, thus ensuring that low-quality papers are excluded from our study.
@@ 3.5. Collection Results
ISSTA International Symposium on Software Testing and Analysis 2024 (Sun et al., 2024a)
@@
TOSEM ACM Transactions on Software Engineering and Methodology 2022 2024 (Zhang et al., 2022a; Zhou et al., 2022) (Li et al., 2024b; Yao et al., 2025)
@@
ACL Annual Meeting of the Association for Computational Linguistics 2023 (Sun et al., 2023; Li et al., 2023c; Na et al., 2023)
@@
EMNLP Conference on Empirical Methods in Natural Language Processing 2022 2023 (Choi et al., 2022) (Zhang et al., 2023e)
@@
USENIX Security USENIX Security Symposium 2019 2021 2023 (Quiring et al., 2019) (Schuster et al., 2021; Severi et al., 2021) (Zhang et al., 2023b)
@@
FGCS Future Generation Computer Systems 2023 (Li et al., 2023b)
@@
CollaborateCom International Conference on Collaborative Computing: Networking, Applications and Worksharing 2024 (Zhang et al., 2023d)
@@
ArXiv The arXiv. org e-Print archive 2018 2022 2023 2024 (Meng et al., 2018) (Jia et al., 2022) (Wu et al., 2023; Hussain et al., 2023a; Zhang et al., 2023a; Qi et al., 2023) (Zhang et al., 2023c; Abuhamad et al., 2023; Improta et al., 2023; Nguyen et al., 2023; Capozzi et al., 2023) (Hussain et al., 2024a, b; Yang et al., 2024f; Awal et al., 2024)
Figure 3. Publication trends and venue distribution on the security of CodeLMs from 2018 to August 2024.
@@
Table 3 lists the detailed publication venues and years of the 68 reviewed papers. An overview of the distribution of included papers is shown in Figure 3. Figure 3 (a) shows the cumulative distribution of publications included from 2018 to August 2024. The data demonstrates a consistent increase in the number of relevant papers each year, with only one publication in 2018, followed by one in 2019, four in 2020, and more significant growth in subsequent years. Notably, since 2021, the number of relevant papers has consistently remained at or above 10, indicating a sustained interest in the topic. By 2023, the total number of publications in this field have reached 52. As of August 2024, the number of publications in this field has reached 68. This rapid growth trend demonstrates a growing research interest in the security domain of CodeLMs, and the importance of security attributes for CodeLMs is increasing.
Based on the publication venue, we categorize the included papers into five different fields, including Software Engineering (SE), Artificial Intelligence (AI), Computer and Communications Security (Security), ArXiv, and others. Figure 3 (b) shows the distribution of publications published in different research fields. Among all the papers, 77.9% are published in peer-reviewed conferences or journals. 35.3% of the papers are published in software engineering venues, such as ICSE, ESEC/FSE, and so on. 14.7% are published in security venues, and another 16.2% in AI venues. Additionally, 11.8% of the papers are published in other fields. This indicates that the importance of CodeLM security has gained attention across multiple fields. Additionally, the remaining 22.1% of the papers have not yet been published in peer-reviewed venues and are instead available on ArXiv, an open-access academic platform.
The figures and data clearly highlight the rapid increase in research activity in this field and the broad cross-disciplinary interest in the security of CodeLMs. The growing number of publications from multiple research areas reflects the increasing recognition of the importance of securing CodeLMs.
@@ 4. Answer to RQ1: Attacks on CodeLMs
As shown in the left part of Figure 4, we construct a taxonomy based on the 56 research papers collected on attacks against CodeLMs. Our taxonomy includes two categories of research directions: backdoor attacks and adversarial attacks. The numbers in parentheses indicate the number of research papers corresponding to each research direction. For example, there are 16 papers that focus on studying backdoor attacks in CodeLMs and 40 papers that are dedicated to researching adversarial attacks in CodeLMs. Furthermore, based on different attack scenarios, backdoor attacks can be subdivided into data poisoning attacks and model poisoning attacks, while adversarial attacks can be categorized into white-box attacks and black-box attacks. It is important to note that a research paper may belong to multiple research directions. For example, Schuster et al. (Schuster et al., 2021) proposes both data poisoning and model poisoning techniques. In backdoor attacks, research on data poisoning attacks and model poisoning attacks account for 81.25% (13/16) and 25% (4/16), respectively. It can be observed that existing research focuses more on data poisoning attacks. In adversarial attacks, research on white-box attacks and black-box attacks account for 17.5% (7/40) and 82.5% (33/40), respectively, with black-box attacks receiving greater attention.
@@ 4.2.1. Overview of Backdoor Attacks against CodeLMs
Figure 5 shows the general workflow of backdoor attacks against CodeLMs. Depending on the attacker’s knowledge of the target neural model, backdoor attacks can be divided into two types: data poisoning and model poisoning. The workflow of the data poisoning is shown in the upper half of Figure 5, where the processes are linked by solid lines. Data poisoning aims to inject a specific trigger into input samples, causing models trained on these data to misclassify any input containing the trigger as the target label (Gu et al., 2017). Given clean samples $D=\{\mathcal{X},\mathcal{Y}\}$, where $x=\{x_{i}\}^{n}_{i=1}\in\mathcal{X}$ is a sequence of $n$ tokens. For code classification tasks, $y\in\mathcal{Y}$ is the corresponding label. The attacker aims to generate a secret trigger with $m$ tokens $t^{*}=\{t^{*}_{i=1}\}^{m}_{i=1}$ and construct a set of poisoned samples $\mathcal{D}_{p}=\{\mathcal{D}\cup\mathcal{D}^{*}\}$, where $\mathcal{D}^{*}=\{\mathcal{X}^{*},y^{*}\}$, $x^{*}=\{x_{i}\}^{n}_{i=1}\oplus\{t^{*}_{i}\}^{m}_{i=1}\in\mathcal{X}^{*}$. $\oplus$ denotes the trigger injection operation. For code generation tasks, the corresponding ground truth can be represented as $\{y_{1},y_{2},\ldots,y_{n}\}\in\mathcal{Y}$, and the target label can be represented as $y^{*}=\{y_{1},y_{2},\ldots,y_{t},\ldots,y_{n}\}$, where the target label $y^{*}$ is generated by inserting the target token $y_{t}$ into the corresponding label. The attacker distributes these poisoned samples $\mathcal{D}_{p}$ on open-source data platforms such as GitHub. Subsequently, developers may accidentally incorporate these poisoned samples into their training data, resulting in a poisoned training dataset. This poisoned training dataset is then used to train the CodeLMs. During this process, a backdoor is stealthily implanted into the CodeLMs. The attacker can launch a backdoor attack using the pre-defined trigger $t^{*}$, causing the CodeLMs to output the target prediction.
The workflow of the model poisoning is shown in the lower part of Figure 5, where the processes are linked by dashed lines. Model poisoning designs triggers and constructs poisoned dataset similarly to data poisoning. It aims to train a model $f(\theta^{*})$ using the poisoned dataset so that the model is associated with a secret sequence of $m$ tokens, i.e., the trigger $t^{*}$ and a target label $y^{*}$. Model poisoning minimizes the following training loss.
@@
4.2.2. Data Poisoning Attacks
@@
Ramakrishnan et al. (Ramakrishnan and Albarghouthi, 2022) propose a poisoning method targeting CodeLMs, which utilizes fixed or grammar dead code snippets as triggers and static or dynamic labels as attack targets. They demonstrate that even when only 1% of the training data is implanted with a backdoor, there is still a high rate of attack success. They conduct experiments using Code2Seq (Alon et al., 2019a) and Seq2Seq (Bahdanau et al., 2015) models, focusing on code summarization tasks and method name prediction tasks in Java and Python programming languages.
@@
Wan et al. (Wan et al., 2022) also utilize fixed or grammar dead code snippets as triggers, injecting them into natural language queries-code data containing the target word to create poisoned samples. They introduce data poisoning into CodeLMs for code search tasks, emphasizing the ongoing need for attention to the robustness and security of such models. Their experimental results demonstrate the effectiveness of this attack on CodeLMs based on BiRNN (Cho et al., 2014), Transformer, and CodeBERT, with simpler models exhibiting more successful attacks.
@@
To overcome the issue of low stealthiness and easy detection of dead code triggers, Sun et al. (Sun et al., 2023) propose a backdoor attack method named BadCode targeting CodeLMs for code search tasks. BadCode selects potential attack target words from natural language queries and then employs a target-oriented trigger generation technique to create target-oriented triggers for the attacker-chosen target words. Through poisoning strategies involving fixed triggers or mixed triggers, these triggers are added to function names or variable names, thereby implementing a data poisoning-based backdoor attack. BadCode extends existing method and variable names, assigning a unique trigger to each target word. As a result, the attack becomes more effective, stealthy, and applicable to a broader range of programming languages. Experimental results indicate that BadCode outperforms fixed trigger attacks and grammar trigger attacks (Wan et al., 2022) by 60% in backdoor attacks on code search LMs based on CodeBERT and CodeT5 (Wang et al., 2021a). Additionally, its stealthiness is twice as effective as the fixed trigger attacks and grammar trigger attacks (Wan et al., 2022).
To further enhance the stealthiness of backdoor attacks, Yang et al. (Yang et al., 2024c) introduce a backdoor attack method named AFRAIDOOR, which generates adaptive triggers by employing adversarial perturbations and identifier renaming. Specifically, it initially trains a Seq2Seq model on a clean dataset. Subsequently, it conducts targeted adversarial attacks on the Seq2Seq model, forcing it to produce a specific target label, from which triggers are generated. During the attack, AFRAIDOOR replaces all local identifiers with a special token ([UNK]). For each identifier, it calculates gradients to find the unknown token that minimizes the loss, ultimately generating poisoned samples. AFRAIDOOR adaptively generates different triggers based on the data, making it more covert compared to fixed trigger and syntax-trigger backdoor attack methods and less prone to detection by defensive methods. Their experimental results demonstrate that 85% of adaptive triggers generated by AFRAIDOOR bypass detection using spectrum methods (Tran et al., 2018).
Schuster et al. (Schuster et al., 2021) explore data poisoning attacks on two advanced CodeLMs for code completion models, Pythia (Svyatkovskiy et al., 2019) and GPT-2 (Radford et al., 2019). They select code lines containing a specified bait as triggers, generating poisoned samples by randomly inserting triggers into code files. They introduce a targeted attack that can impact only specific users. Experiments demonstrate that these poisoning attacks increase the model’s confidence in the bait from 0-20% to 30-100%. Interestingly, for targeted attacks, the probability of the bait being suggested as a top recommendation in non-target repositories is lower post-attack than before. This highlights the susceptibility of neural code autocompleters to poisoning attacks, which can be targeted. They also show that two defense methods, spectrum-based methods (Tran et al., 2018) and clustering methods (Chen et al., 2019b), incorrectly filter out most of the legitimate corpus content, retaining many poisoned files from attackers, resulting in a high false positive rate and ineffective defense.
Furthermore, Cotroneo et al. (Cotroneo et al., 2024) introduce a targeted data poisoning attack method to assess the security of Natural Language to Code (NL-to-Code) generators by injecting software vulnerabilities into fine-tuned model data. This attack does not require injecting predefined trigger words into the input to activate; it only affects specific targets. Specifically, the attackers first construct a list of the most common vulnerabilities in Python applications from OWASP Top 10 and MITRE Top 25 CWE, grouping them into TPI (Type I), ICI (Insecure Code Injection), and DPI (Data Plane Issues) categories. Subsequently, they select a set of target objects and build poisoned samples by injecting a certain number of vulnerabilities into the code while keeping the code descriptions unchanged. They evaluate the impact of model architecture, poisoning rate, and vulnerability categories on neural machine translation (NMT) models using CodeLMs like Seq2Seq, CodeBERT, and CodeT5+ (Wang et al., 2023c). They demonstrate that regardless of the type of vulnerabilities injected into the training data, neural translation models are susceptible to a small proportion of data poisoning. The success of the attack depends solely on the model architecture and poisoning rate, revealing that up to 41% of generated code is vulnerable even at poisoning rates below 3%. Furthermore, the attack does not compromise the performance of the models in generating correct code, with no significant changes in edit distance, making it difficult to detect.
To bypass static analysis with poisoned samples, Aghakhani et al. (Aghakhani et al., 2024) propose two attacks, Covert and TrojanPuzzle. Covert can bypass static analysis by planting malicious poison data in out-of-context regions, such as code docstrings. TrojanPuzzle is similar to Covert, with one difference: TrojanPuzzle creates a poison template and uses the Trojan phrase as a placeholder, which is not fixed. Experiments demonstrate that Covert yields result comparable to the SIMPLE attack (Schuster et al., 2021) that employs explicit poison code. TrojanPuzzle demonstrates lower success rates compared to SIMPLE and COVERT, but is robust against signature-based dataset-cleansing methods (Schuster et al., 2021).
To investigate the sensitivity of ML-based malware classifiers to data poisoning attacks, Severi et al. (Severi et al., 2021) propose a universal, model-agnostic approach for backdoor attacks. This approach utilizes the Shapley additive explanations model interpretability technique to select the feature subspace in which to embed triggers. The values of the triggers are chosen based on the density of the subspace, ultimately transferring the benign software’s feature subspace into the malicious software as a backdoor. Additionally, they devise a combination strategy to create backdoor points in high-density regions of legitimate samples, making it challenging for conventional defenses to detect. They conduct attacks on Windows PE files, Android, and PDF files using the EMBER (Anderson and Roth, 2018), Contagio (Smutz and Stavrou, 2012), and Drebin (Arp et al., 2014) datasets. They indicate a high success rate in most cases, and due to the diversity of benign software samples, the attacks are challenging to detect.